the mass of water in grams formed by the reaction when 11.2 L of H2 reacts with 5.6 L of O2

Grade:12th pass

3 Answers

Arun
5 years ago
Dear student
 
H2 + ½ O2 --------> H2O
 
hence 11.2 mole water forms or ½ mole
 
hence
 
½ = w / 18
 
w = 9 gm
 
hpe it helps
 
Khimraj
5 years ago
22.4 L = 1 mole 
so mole of H2 = ½ 
moe of O2 =  1/4
then complete H2 and O2 react to form ¼ mole of water
so mass of water = 18/4 = 4.5 gm
Hope it clears.
Manda Snehashish Reddy
5 years ago
We very well know the reaction 2H2 + O2 = 2H2O. It is balanced.
We can see that 2 moles of H2 and 1 mole of O2 combine to form 2 moles of water.
Let us convert the given information into moles.
11.2 Litres is 0.5 moles
5.6 Litres is 0.25 moles.
According to reaction 2 + 1 gives 2
We have,               0.5 + 0.25 which must give 0.5 moles of H2O.
 
Thus, the answer is 9 grams.
                                          Regards, M.S.R